# Blockchain with proof of quantum work

### v2

@@ C.2.1 Partial Quantum Work Attack
Since a finite fraction of stakeholders (thence mining power) can validate a near-threshold block, there is finite probability of such a block being incorporated in the strongest chain subject to a delay; motivating a minor to broadcast such a block. A careful balancing of block penalties and rewards is necessary to ensure quantum-partial work is not a profitable strategy, achievable by control of transaction and block rewards and enhanced block structure.
The quantum experimental set-up also allows room for efficiencies as well, which should be understood in a full application. To minimize QPU access time (costs) a miner should optimize parameters such as number of samples and programming time. They could then more efficiently search, or filter, nonces. A miner is incentivized to prefilter nonces with cheap experiments, and then seek higher confidence (via enhanced sampling) only on candidates for broadcast. In our proof-of-concept demonstrations by contrast we fix QPU access time to 1 second in mining and validation.
@@ C.3 Multi-Block Attack
Our definitions of Chainwork mean that for any path to be viable as a strongest chain there must be strictly more verified blocks than unverified blocks. Owing to this, the criteria necessary for this attack to succeed are similar to the standard 51% attack. In effect an attacker must be able to grow the chain more quickly than alternative non-fraudulent candidates, in order to undo the Chainwork damage of their fraudulent block(s). This requires control of a significant fraction of community compute power assuming practical chain efficiency (concentrated honest mining power, up to small fluctuations).
@@ C.4 Enhanced Block Structure Mitigations
An enhanced block structure shown in Fig. 6 can mitigate against no-work and partial-work attacks, as now discussed.
Figure 6: The standard block (details in Figure 1) can be broadcast with additional information or costs, shown in red and blue, to mitigate for weaknesses. Some or none of these enhancements may be useful, depending on the strength of the quantum work. Some of this information, marked red, such as the enhanced experimental data, is not cryptographically protected by the quantum work, but this encoding becomes cryptographically (securely) linked when a subsequent block is created, via the SHA-256 hash of the enhanced block (entering the next standard header). Hybrid (classical) work (characterized by a secondary nonce requirement), or stakes, can be a means to raise the cost of tampering with this information, and/or the costs associated with fraudulent block broadcasts. A stake is a signed payment by the miner that is forfeited if the block fails to enter the strongest chain (not related to proof of stake). Stakes associated with failed branching can be collected alongside the standard transactions, disincentivizing branch proposal. The secondary block information may be marked with fields mirroring the standard block, such as timestamp, version and hardness (not shown).
Block broadcasts that are verified with low probability pay a broadcasting cost. A sufficient strategy for mitigation is raising the cost of broadcasting blocks that have low-probability of validation, and raising the bar for validation by requiring the publication of enhanced experimental information. A modified block structure can achieve this as shown in Fig. 6. We describe some potential components:
C.4.1 Network and identity mitigations
If broadcasters are not anonymous their behavior can be tracked. A rule can be introduced to penalize or temporarily block users who consistently propose unverifiable work, as is done for DoS. Note that a transmission path or node might be penalized with limited information on the stakeholder.
C.4.2 Broadcasting with additional classical work
@@
In order to broadcast it can be required that a user completes a classical proof of work. The work should be enough to delay transmission and raise costs, yet negligible compared to the quantum work. If the classical work is sufficiently high, it ceases to become profitable to risk approximations in the quantum part. In the case where the classical work becomes comparable to the quantum work, we have a hybrid system. In an immature technology it may be valuable to pursue such a hybrid approach and gradually reduce the proportion of classical work as confidence is gained in the quantum operation. Quantum proof of work could of course also be hybridized with proof of stake as an alternative mitigation.
C.4.3 Broadcasting with a stake
@@
A user might be required to stake (bid on) their proposal. The stake can be a finite fraction of the block reward, signed from the miner’s current wallet to a burn wallet. Stakes for failed branches are paid to a burn-wallet. In this way, any user is limited in the number of weak proposals they are able to make, and risk significant loses from doing so. Payment to a burn wallet is a standard component in proof of stake blockchains such as Etherium, but note that the stake here does not require any community involvement. This is not a proof-of-stake consensus mechanism, which requires a community-wide mechanism.
C.4.4 Enhanced experimental data
We can also raise the difficulty to pass validation by requiring the publication of enhanced experimental data (EED). Quantum work is based on a $\mathcal{N}_{\rm zeros}$ bit description of the experimental outcome, by contrast the full experimental data may be subject to more robust statistical validation tests that more robustly eliminate bad proposals [1]. EED can fall outside the header if the information content is large, and be linked in a similar manner to transaction data. The $\mathcal{N}_{\rm zeros}$ bits used in validation are cryptographically protected, but a user might be required to publish EED consisting of more extensive statistics (up to perhaps full sample sets) as part of a broadcast. Although EED might be manipulated, it might be taken as a rule to evaluate the strongest available EED. Weak classical work (2.) or a stake (3.) might be used to protect EED and disincentivize manipulation. The EED can be hashed on the next mining event, so that it is cyrptographically protected subject to a delay of one block. For quantum-supremacy satisfying experiments EED might not be spoofed by classical methods, but this does not prevent classical filtering of nonces as described above.
@@ C.5 Hyperplane distribution mitigations
Figure 7: Operation of an example blockchain of 128 block broadcasts generated in the limit of many miners using projection orthogonal to $J$ with $\mathcal{N}_{zeros}=32$, and confidence-based Chainwork with $\mathcal{N}_{\rm zeros}=1$. The most recent block mined is marked in black and the strongest chain in blue. Averaging efficiency across 5 independent chains yields an efficiency estimate of 0.34 +/- 0.03 in aggrement with the bootstrapped estimates of Figure 16.
Suppose a per-nonce classical work less than the quantum experimental cost ($C_{H}$) allows for a witness estimator ${\hat{W}}$. One mitigation is to modify the distribution of thresholds or hyperplanes (to choose $W_{0}\sim{\hat{W}}$), or to choose hyperplanes orthogonal to ${\hat{W}}$. In either case the classical approximation is rendered uninformative on the signs (bits) expected from the quantum experiment. If the uncertainty in the quantum experiment is relatively small, we might still have sufficient accuracy in quantum experiments for robust validation. A visualization of this process is shown in Fig. 12.
In the case of our quench experiments, we know that nearest-neighbor correlation are correlated with $-J$, the approximation $C_{ij}\propto-J_{ij}$ describes accurately the weak coupling (fast anneal) limit. There are significant deviations from this estimate, but any hyperplane that is well correlated with $J$ can accumulate a reliable signal from weak estimators incorporating this approximation. To mitigate for this we might choose our hyperplanes to be orthogonal to $J$. Choosing hyperplanes (or thresholds $W_{0\alpha}$) to decrease reliability of classical estimators will typically enhance the uncertainty in quantum validation, thus reducing the efficiency and enhance delays in a blockchain. This effect is shown in Fig. 16 as a function of $\mathcal{N}_{zeros}$. In Fig. 7 we show an example of a chain resistant to simple classical inference based on the weak coupling limit approximation.
@@ Appendix D Experimental Implementation Details
Figure 8: Energy scales (21) of the general-access QPUs used in experiments. $\Gamma(s)$ decreases with anneal progression $s$, ${\cal J}(s)$ increases. Transverse fields are comparable on all devices, but the problem energy scale is larger in Advantage2. Advantage QPUs can be programmed with (approximately) doubling the anneal duration to simulate an Advantage2 QPU full-energy scale anneal as shown in Figure 9 [1].
@@
Figure 10: Operation of an example blockchain using four generally-accessible Advantage and Advantage2 QPUs with $\mathcal{N}_{\rm zeros}=64$ and 100 miners with data collected across more than 2 weeks, with 1041 total block broadcasts and using basic +/-1 Chainwork. The color scheme matches that of Figure 3. Node placement and line-thickness conventions follow Figure 4. It is apparent in the larger number of black and gray (unconfirmed) transactions, and in the increased rate of edges spanning the spiral, that this blockchain is of lower efficiency and larger delay than that presented in Figure 4. This is consistent with larger $\mathcal{N}_{\rm zeros}$ as shown in Figure 5 and 16. Where a solver became unavailable mining and verification was redistributed on available solvers; the only significant disruption of four-solver availability was a four-hour outage on April 15th owing to a severed internet cable at the Information Sciences Institute in California (Advantage_system6.4). Although the cross validation rate fluctuated through such events, this had only a small impact on the stability of the blockchain and the measured efficiency is well matched to the bootstrapped prediction of Figure 5 at 53 %.
@@
Figure 11: Operation of an example blockchain using the Advantage2 QPU and the dimerized biclique spin-glass problems with $\mathcal{N}_{\rm zeros}=64$ and 100 miners, 1101 block broadcasts and using basic +/-1 Chainwork. Like Figure 10, the efficiency is estimated at 53%, with data collected across more than 2 weeks. Despite differences in parameterization of the cubic and biclique experiments the median cross-validation rate is comparable, consistent with the matched efficiency. The color scheme, node placement and line thicknesses match the conventions of Figures 3 and 4
@@
Figure 12: A high-dimensional statistics vector $V$ is estimated; here only 2 dimensions are shown. Quantum (and classical) estimators are subject to some uncertainty represented as a ball, in high dimensions and with reasonable variance in estimation, all estimators become well resolved in space. By standard locality-sensitive hashing we can digitalize the result. We record what side of each hyperplane the statistics are on. With respect to hyperplanes $G_{2}$ and $G_{3}$ quantum statistics are encoded as +1 and +1, but the outcome with respect to $G_{3}$ is subject to experimental uncertainty. A weak classical estimator, such as any estimator without access to $J$, produces uniform random results. Stronger estimators might produce outcomes correlated in space with the quantum result. Hyperplane distributions can be adjusted by displacement of the origin, or selecting hyperplanes orthogonal to strong classical estimators. $G_{2,\cdot}$ (a projection orthogonal to the strong classical estimator) allows high confidence in the quantum result, but both the strong and weak classical estimators shown produce outcomes uncorrelated with the quantum result.
@@
Figure 13: Validation rates with respect to experiments on different QPUs. Validation rates are higher using a common QPU, or a common QPU architecture (Advantage (Adv) or Advantage2 prototype (Adv2)). They are slightly higher on Advantage QPUs due the smaller size of the prototype Advantage2 prototype QPU combined with the bound of 1 second of QPU access time. On each QPU the probability of an outcome 1 is measured on each QPU from $32$ independent (with randomized embeddings) programmings, for 20 $J$ realizations and total of $20480$ projections. From these $4\times 20\times 32$ programmings we infer bit error rates for different QPU combinations. Three Advantage QPUs are used, the relatively small variability between these QPUs is not shown. (top) Hyperplane are chosen with independent and normally distributed components. (bottom) With the same hyperplanes projected into the space orthogonal to $J$ (see Section C.5).
@@
Figure 14: Witnesses $W_{\alpha}$ measured on Advantage (Adv) and Advantage2 (Adv2) QPUs are similar, and much larger than the root mean square (RMS) difference between $W_{\alpha}$ in different experiments (programmings with independent embeddings on a fixed system, or between systems, as labeled). $\delta W_{\alpha}$ is shown as a root mean square value across all programmings (variation of embedding and QPUs). (top) Each hyperplane has independent and normally distributed components. (bottom) With the same hyperplanes projected into the space orthogonal to $J$. This enhances spoof-resistance, but reduces the signal to noise ratio ($W_{\alpha}$ are made smaller relative to $dW_{\alpha}$). (top) Hyperplane are chosen with independent and normally distributed components. (bottom) With the same hyperplanes projected into the space orthogonal to $J$ (see Section C.5).
@@
Figure 15: Expected hashes required per broadcast with confidence-based Chainwork measured by Monte Carlo sampling. For basic +/- Chainwork the expected number of hashes required per broadcast is precisely $2^{\mathcal{N}_{\rm zeros}}$; confidence-based Chainwork is close to this, with corrections shown. As $\mathcal{N}_{\rm max}$ increases blocks are easier to mine. As $\mathcal{N}_{\rm max}$ increases it is easier to find work-satisfying hashes. Using J-orthogonal hyperplanes more bits are uncertain and it becomes harder to find a confidence-satisfying bit sequence (mining rate goes down).
@@
Figure 16: Statistics from 16 chains, each of length 512-1024 using the QPU-witness bootstrapping methodology (see Appendices D.1 and D.3). Hyperplanes are either i.i.d normally distributed, or are restricted to be orthogonal-to-$J$ (strengthening hashes against weak classical estimators). Resistance to spoofing increases uncertainty in validation, reducing efficiency. The Chainwork uses either the basic +/-1 definition, or the confidence-weighted definition with $\mathcal{N}_{\rm max}=1$ or $2$: confidence-weighted strongest chains are more efficient. Statistics are established via the bootstrapping method, with statistic in agreement with the full experiments shown in Figures 4, 7 and 10. (top) Mining efficiency measures the fraction of block broadcasts entering the strongest chain in the limit of many miners. (bottom) Transactions (blocks) are uncertain until agreed upon by the majority of miners. Transaction delay with 95% confidence measures the depth required for a consolidation of 95% of mining power in the limit of many miners. Stakeholders might consider blocks buried this deep within their strongest chain as immutable with 95% confidence; higher confidence implies higher delays. Data is truncated where mining efficiency drops below 0.2, as mining depth becomes strongly biased as a function of the finite blockchain size.
@@
The dimerized cubic lattice is equivalent to a cubic lattice minor-embedded in the QPU with chain strength 1, with a regular pattern of connectivity between chains. All couplers are ±$\pm 1$ uniformly and independently distributed. This matches the cubic-nodimer ensemble of [1]. A dimerized biclique is equivalent to an embedded biclique ($K_{n,n}$) model. Couplers are ±$\pm 1$ on chains, and ±$\pm\frac{1}{\sqrt{n}}$ between chains in a regular pattern. For the dimerized biclique, the regular pattern of inter-chain couplers is specific to the Advantage2 QPU, so we do not perform mining or cross-QPU validation on Advantage QPUs. The dimerized bicliques ensemble matches that of [1].
[…]
@@
Per experiment (mining or validation) we choose uniformly at random among the available QPUs and then embed the problem subject to an automorphism and spin-reversal transform on each available embedding. The effect of these transformations is to randomize the mapped position and signs on programmed fields and couplers, in effect emulating the enhanced noise that might be anticipated from (more) distributed computation.
@@
Returned sample sets are processed to nearest-neighbor correlations $C$ (1 per programmed coupler), that are then projected with independent and identically distributed hyperplanes to determine witnesses. Hyperplanes have independent and normally distributed components. In the case of Fig. 16 we also consider a variation in which hyperplanes are projected into the space orthogonal to $J$, as described in C.5. $C$ is correlated with $-J$, and we remove the linear part of this correlation with such a projection, making the quantum statistics harder to spoof classically. A visualization of the projection process is shown in Fig. 12, experimental distributions for witnesses are shown in Fig. 14. Uncertainty in witness values is typically much smaller than the witness value. Witness values vary more between Advantage and Advantage2 QPUs, than for same-generation QPU validations.
@@
Witnesses are thresholded to determine hashes. This digitalization of the correlations is the standard locality-sensitive hashing method, which guarantees preservation of local distances. With enough projections a small (supremacy achieving error) in correlations translates to a small (supremacy-achieving) Hamming distance in the resulting bit sequence. The probability that bits are validated (reproduced) under a repeated experiment are shown in Fig. 13. More than 99% of bits are typically validated, but the remaining bits can have high uncertainty, particularly with respect to cross-processor values.
@@
We performed simulations of many chains with 100 miners modeling a mining pool and transactions, or taking the limit of many miners (where no miner mines more than one block). Examples are shown in Figs. 4, 10, 11 and 7. However, only the distribution of witness statistics impacts the structure and stability of the chain, assuming no simultaneous broadcasts, via mining events and stakeholder validation. Understanding this we were able to accelerate mining (Section D.2), and simulate experiments by bootstrapping (Section D.3). We first describe our main statistical measures of blockchain stability in Section D.1.

### v3

@@ C.2.1 Partial Quantum Work Attack
Since a finite fraction of stakeholders (thence mining power) can validate a near-threshold block, there is finite probability of such a block being incorporated in the strongest chain subject to a delay; motivating a miner to broadcast such a block. A careful balancing of block penalties and rewards is necessary to ensure quantum-partial work is not a profitable strategy, achievable by control of transaction and block rewards and enhanced block structure.
The quantum computation also allows room for efficiencies, which should be understood in a full application. To minimize QPU access time (costs) a miner should optimize parameters such as number of samples and programming time. They could then more efficiently search, or filter, nonces. A miner is incentivized to prefilter nonces with cheap experiments, and then seek higher confidence (via enhanced sampling) only on candidates for broadcast. In our proof-of-concept demonstrations by contrast we fix QPU access time to 1 second in mining and validation.
@@ C.3 Multi-Block Attack
Our definitions of Chainwork mean that for any path to be viable as a strongest chain there must be strictly more verified blocks than unverified blocks. Owing to this, the criteria necessary for this attack to succeed are similar to the standard 51% attack. In effect an attacker must be able to grow the chain more quickly than alternative non-fraudulent candidates, in order to undo the Chainwork damage of their fraudulent block(s). Although a user might time an attack to coincide with a moment when compute power is fragmented, this still requires control of a significant fraction of community compute power assuming practical chain efficiency (concentrated honest mining power, up to small fluctuations).
@@ C.4 Enhanced Block Structure Mitigations
An enhanced block structure shown in Fig. 6 can mitigate against several of the attacks discussed. The new block consists of fields known prior to quantum-hash estimation, which impact the parameterization of the quantum-hash estimator, and post-quantum-hash fields, for which manipulation is prevented by the miner’s cryptographic signature.
Figure 6: The standard block (details in Figure 1) can be broadcast with additional information or costs to mitigate for weaknesses. Some or none of these enhancements may be useful, depending on the strength of the quantum hash. Dark green and light green are elements that are known prior to the quantum-hash estimation, light being new elements discussed in this appendix. Additional elements in blue can be broadcast and signed by the miner. Since the public key of the miner is protected by completion of the work, it is not possible for anyone but the miner to manipulate bits of the enhanced broadcast without repeating the proof of work. Completion of supplementary classical work, or inclusion of transactions costs can raise the cost of tampering with this information, and/or the costs associated with fraudulent block broadcasts. One may consider the work requirement with respect to the quantum hash, or a cryptographic hash of the enhanced header. In the latter case, this allows a cheap classical check of block validity, before a second more expensive quantum computation stage must be undertaken.
C.4.1 Broadcasting with Additional Classical Work
@@
Block broadcasts that are verified with low probability pay a broadcasting cost. In order to raise the cost of broadcasting it can be required that a user completes a classical proof of work. The work should raise costs, yet be negligible compared to the quantum work. If the classical work is sufficiently high, it ceases to become profitable to risk approximations in the quantum part. In the case where the classical work becomes comparable to the quantum work, we have a hybrid system. In an immature technology it may be valuable to pursue such a hybrid approach and gradually reduce the proportion of classical work as confidence is gained in the quantum operation. Quantum proof of work could of course also be hybridized with proof of stake as an alternative mitigation.
C.4.2 Broadcasting with a Stake
@@
The cost of broadcasting is also raised if a user is required to stake (bid on) their proposal. The stake can be a finite fraction of the block reward, signed from the miner’s current wallet to a burn wallet. Stakes for failed branches are paid to a burn-wallet. In this way, any user is limited in the number of weak proposals they are able to make, and risk significant loses from doing so. Payment to a burn wallet is a standard component in proof-of-stake blockchains such as Etherium, but note that the stake here does not require any community involvement. This is not a proof-of-stake consensus mechanism, which requires a community-wide mechanism. Stake payments from blocks in branches that were ultimately rejected by consensus can later be collected by miners extending the strongest chain, so that they become part of the permanent ledger record. Thus even when a block is rejected, its stake must still be paid; this disincentivizes the broadcast of poor-quality blocks.
C.4.3 Enhanced Experimental Data
We can also raise the difficulty to pass validation by requiring the publication of enhanced experimental data (EED). Quantum work is established based on the leading $\mathcal{N}_{\rm zeros}$ bit description of the quantum hash; by contrast the high-dimensional full experimental data may be subject to more-robust statistical validation tests that detect bad proposals [1]. Assuming we choose our unitary evolution ensemble so that a plurality of quantum unitary evolutions are not simulatable (to accuracy comparable to that of the quantum-computing participants) classical computing can be eliminated with high probability by choice of a distance threshold (in statistic space, per [1]). EED can be signed by the miner’s private key matching the public key that is protected by work. Only the miner can manipulate the experimental data without redoing the work, and is incentivized to demonstrate the best (true) data.
C.4.4 Enhanced-block Hash for Work Mitigation
In place of evaluating $N_{zeros}$ on the quantum hash, we may append a quantum hash to the pre-experimental block header and require that a classical cryptographic hash of this object meet the $N_{zeros}$ requirement. This hash can also be used as the block identity, in place of the hash based only on the pre-experimental header components.
This scheme allows for a cheap classical filter that quickly detects manipulation: Before conducting a quantum experiment, or validating other block fields, we can hash the enhanced block header to check it meets the proof-of-work threshold ($N_{zeros}$). This is by contrast with the alternative scheme in which a quantum experiment is required to check the work. Note that this test allows for false positives: it is necessary to check the experimental data via a statistical test to ensure a capable quantum computer actually ran the experiment.
The quantum hash must be consistent with the extended experimental data, must be long enough that a verifier has confidence the broadcaster undertook the valid quantum experiment, yet short enough that miners cannot confidently manipulate the experimental data to attempt many proofs of work from a single quantum experiment (per partial quantum work section C.2.1). Note that the length of the quantum hash needn’t be determined by ${\mathcal{N}}_{zeros}$, as is the case in our simplest implementation.
@@ C.5 Hyperplane Distribution Mitigations
Figure 7: Operation of an example blockchain of 128 block broadcasts generated in the limit of many miners using projection orthogonal to $J$ with $\mathcal{N}_{zeros}=32$, and confidence-based Chainwork with $\mathcal{N}_{\rm zeros}=1$. The most recent block mined is marked in black and the strongest chain in blue. Averaging efficiency across 5 independent chains yields an efficiency estimate of 0.34 +/- 0.03 in agreement with the bootstrapped estimates of Figure 16.
Suppose a per-nonce classical work less than the quantum experimental cost ($C_{H}$) allows for a witness estimator $^{\hat{W}}$. One mitigation is to modify the distribution of thresholds or hyperplanes (to choose $W_{0}\sim{\hat{W}}$), or to choose hyperplanes orthogonal to $^{\hat{W}}$. In either case the classical approximation is rendered uninformative on the signs (bits) expected from the quantum experiment. If the uncertainty in the quantum experiment is relatively small, we might still have sufficient accuracy in quantum experiments for robust validation. A visualization of this process is shown in Fig. 12.
In the case of our quench unitary dynamics, we know that nearest-neighbor correlation are correlated with J-J: the approximation $C_{ij}\propto-J_{ij}$ describes accurately the weak coupling (fast anneal) limit. There are significant deviations from this estimate, but any hyperplane that is well correlated with $J$ can accumulate a reliable signal from weak estimators incorporating this approximation. To mitigate for this we might choose our hyperplanes to be orthogonal to $J$. Choosing hyperplanes (or thresholds $W_{0\alpha}$) to decrease reliability of classical estimators will typically enhance the uncertainty in quantum validation, thus reducing the efficiency and increasing delays in a blockchain. This effect is shown in Fig. 16 as a function of $\mathcal{N}_{zeros}$. In Fig. 7 we show an example of a chain resistant to simple classical inference based on the weak coupling limit approximation.
@@ Appendix D Experimental Implementation Details
Figure 8: Annealing schedule energy scales of the general-access QPUs used in experiments. $\Gamma(s)$ decreases with anneal progression $s$, ${\cal J}(s)$ increases, to define the time-dependent Hamiltonian (21). Transverse fields are comparable on all devices, but the problem energy scale is larger in Advantage2. Advantage QPUs can be programmed with (approximately) doubling the anneal duration to simulate an Advantage2 QPU full-energy scale anneal as shown in Figure 9 [1].
@@
Figure 10: Operation of an example blockchain using four generally-available Advantage and Advantage2 QPUs with $\mathcal{N}_{\rm zeros}=64$ and 100 miners with data collected across more than 2 weeks, with 1041 total block broadcasts and using basic +/-1 Chainwork. The color scheme matches that of Figure 3. Node placement and line-thickness conventions follow Figure 4. It is apparent in the larger number of black and gray (unconfirmed) transactions, and in the increased rate of edges spanning the spiral, that this blockchain is of lower efficiency and larger delay than that presented in Figure 4. This is consistent with larger $\mathcal{N}_{\rm zeros}$ as shown in Figure 5 and 16. Where a solver became unavailable mining and verification was redistributed on available solvers; the only significant disruption of four-solver availability was a four-hour outage on April 15th owing to a severed internet cable at the Information Sciences Institute in California (Advantage_system6.4). Although the cross validation rate fluctuated through such events, this had only a small impact on the stability of the blockchain and the measured efficiency is well matched to the bootstrapped prediction of Figure 5 at 53 %.
@@
Figure 11: Operation of an example blockchain using the Advantage2 QPU and the dimerized biclique spin-glass problems with $\mathcal{N}_{\rm zeros}=64$ and 100 miners, 1101 block broadcasts and using basic +/-1 Chainwork. Like Figure 10, the efficiency is estimated at 53%, with data collected across more than 2 weeks. Despite differences in parameterization of the cubic and biclique demonstrations the median cross-validation rate is comparable, consistent with the matched efficiency. The color scheme, node placement and line thicknesses match the conventions of Figures 3 and 4
@@
Figure 12: A high-dimensional statistics vector $V$ is estimated; here only 2 dimensions are shown. Quantum (and classical) estimators are subject to some uncertainty represented as a ball, in high dimensions and with reasonable variance in estimation, all estimators become well resolved in space. By standard locality-sensitive hashing we can digitalize the result. We record what side of each hyperplane the statistics are on. With respect to hyperplanes $G_{2}$ and $G_{3}$ quantum statistics are encoded as +1 and +1, but the outcome with respect to $G_{3}$ is subject to experimental uncertainty. A weak classical estimator, such as any estimator without access to $J$, produces uniform random results. Stronger estimators might produce outcomes correlated in space with the quantum result. Hyperplane distributions can be adjusted by displacement of the origin, or selecting hyperplanes orthogonal to strong classical estimators. $G_{2}$ (a projection orthogonal to the strong classical estimator) allows high confidence in the quantum result, but both the strong and weak classical estimators shown produce outcomes uncorrelated with the quantum result.
@@
Figure 13: Validation rates with respect to experiments on different QPUs. Validation rates are higher using a common QPU, or a common QPU architecture (Advantage (Adv) or Advantage2 prototype (Adv2)). They are slightly higher on Advantage QPUs due the smaller size of the prototype Advantage2 prototype QPU combined with the bound of 1 second of QPU access time. On each QPU the probability of an outcome 1 is measured on each QPU from $32$ independent (with randomized embeddings) programmings, for 20 $J$ realizations and total of $20480$ projections. From these $4\times 20\times 32$ programmings we infer bit error rates for different QPU combinations. Three Advantage QPUs are used, the relatively small variability between these QPUs is not shown. (Top) Hyperplanes are chosen with independent and normally distributed components. (Bottom) With the same hyperplanes projected into the space orthogonal to $J$ (see Section C.5).
@@
Figure 14: Witnesses $W_{\alpha}$ measured on Advantage (Adv) and Advantage2 (Adv2) QPUs are similar, and much larger than the root mean square (RMS) difference between $W_{\alpha}$ in different quantum-hash calculations (programmings with independent embeddings on a fixed system, or between systems, as labelled). $\delta W_{\alpha}$ is shown as a root mean square value across all programmings (variation of embedding and QPUs). (Top) Each hyperplane has independent and normally distributed components. (Bottom) With the same hyperplanes projected into the space orthogonal to $J$. This enhances spoof-resistance, but reduces the signal to noise ratio ($W_{\alpha}$ are made smaller relative to $dW_{\alpha}$). (Top) Hyperplane are chosen with independent and normally distributed components. (Bottom) With the same hyperplanes projected into the space orthogonal to $J$ (see Section C.5).
@@
Figure 15: Expected hashes required per broadcast with confidence-based Chainwork measured by Monte Carlo sampling. For basic +/- Chainwork the expected number of hashes required per broadcast is precisely $2^{\mathcal{N}_{\rm zeros}}$; confidence-based Chainwork is close to this, with corrections shown. As $\mathcal{N}_{\rm max}$ increases, blocks are easier to mine. As $\mathcal{N}_{\rm max}$ increases, it is easier to find work-satisfying hashes. Using J-orthogonal hyperplanes, more bits are uncertain and it becomes harder to find a confidence-satisfying bit sequence (mining rate goes down).
@@
Figure 16: Statistics from 16 chains, each of length 512-1024 using the QPU-witness bootstrapping methodology (see Appendices D.1 and D.3). Hyperplanes are either i.i.d normally distributed or are restricted to be orthogonal to $J$ (strengthening hashes against weak classical estimators). Resistance to spoofing increases uncertainty in validation, reducing efficiency. The Chainwork uses either the basic +/-1 definition or the confidence-weighted definition with $\mathcal{N}_{\rm max}=1$ or 22: confidence-weighted strongest chains are more efficient. Statistics are established via the bootstrapping method, with statistic in agreement with the full experiments shown in Figures 4, 7 and 10. (Top) Mining efficiency measures the fraction of block broadcasts entering the strongest chain in the limit of many miners. (Bottom) Transactions (blocks) are uncertain until agreed upon by the majority of miners. Transaction delay with 95% confidence measures the depth required for a consolidation of 95% of mining power in the limit of many miners. Stakeholders might consider blocks buried this deep within their strongest chain as immutable with 95% confidence; higher confidence implies higher delays. Data is truncated where mining efficiency drops below 0.2, as mining depth becomes strongly biased as a function of the finite blockchain size.
@@
The dimerized cubic lattice is equivalent to a cubic lattice minor-embedded in the QPU with chain strength 1, with a regular pattern of connectivity between chains. All couplers are $\pm 1$ uniformly and independently distributed. This matches the cubic-nodimer ensemble of [1]. A dimerized biclique is equivalent to an embedded biclique ($K_{n,n}$) model. Couplers are $\pm 1$ on chains, and $\pm\frac{1}{\sqrt{n}}$ between chains in a regular pattern. For the dimerized biclique, the regular pattern of inter-chain couplers is specific to the Advantage2 QPU, so we do not perform mining or cross-QPU validation on Advantage QPUs. The dimerized bicliques ensemble matches that of [1].
For 3D spin-glasses quantum-hash computation (for mining or validation stages) on each of the four generally-available QPUs, in each case with parallel embeddings (because we consider relatively small problems of 128 qubits, we can program in parallel many copies). We fix the annealing time on the Advantage2 device to be $5\text{\,}\mathrm{n}\mathrm{s}$, and fix annealing times for the other three Advantage devices close to $10\text{\,}\mathrm{n}\mathrm{s}$, so as to approximately maximize the cross-platform validation rate (a median with respect to a test set of spin-glass disorder realizations and hyperplanes). Figure 9 shows rescaling factors necessary for a collapse of the ensemble-average energy. These ratios approximately minimize the correlation error, and maximize the validation rate. The difference in annealing times reflects a difference in energy scales between devices. The unitless unitary evolution then becomes well matched, at least throughout the critical region [1]. Energy scales are shown with respect to the Hamiltonian (21) in Fig. 8. We then fix the number of reads to fully utilize the maximum QPU access time ($1\text{\,}\mathrm{s}\mathrm{e}\mathrm{c}\mathrm{o}\mathrm{n}\mathrm{d}$) per programming, and otherwise use default parameters. For the biclique problem we have 72 qubits and fix the Advantage2 annealing time to $15\text{\,}\mathrm{n}\mathrm{s}$. Demonstrations plots relate to the cubic lattice ensemble, unless stated otherwise.
@@
Per quantum hash computation (mining or validation) we choose uniformly at random among the available QPUs and then embed the problem subject to an automorphism and spin-reversal transform on each available embedding. The effect of these transformations is to randomize the mapped position and signs on programmed fields and couplers, in effect emulating the enhanced noise that might be anticipated from (more) distributed computation.
@@
Returned sample sets are processed to nearest-neighbor correlations $C$ (1 per programmed coupler), that are then projected with independent and identically distributed hyperplanes to determine witnesses. Hyperplanes have independent and normally distributed components. In the case of Fig. 16 we also consider a variation in which hyperplanes are projected into the space orthogonal to $J$, as described in C.5. $C$ is correlated with J-J, and we remove the linear part of this correlation with such a projection, making the quantum statistics harder to spoof classically. A visualization of the projection process is shown in Fig. 12, distributions for witnesses are shown in Fig. 14. Uncertainty in witness values is typically much smaller than the witness value. Witness values vary more between Advantage and Advantage2 QPUs, than for same-generation QPU validations.
@@
Witnesses are thresholded to determine hashes. This digitalization of the correlations is the standard locality-sensitive hashing method, which guarantees preservation of local distances. With enough projections a beyond-classical (small) error in correlations translates to a small Hamming distance in the resulting bit sequence. The probability that bits are validated (reproduced) under a repeated experiment are shown in Fig. 13. More than 99% of bits are typically validated, but the remaining bits can have high uncertainty, particularly with respect to cross-processor values.
@@
We performed simulations of many chains with 100 miners modelling a mining pool and transactions, or taking the limit of many miners (where no miner mines more than one block). Examples are shown in Figs. 4, 10, 11 and 7. However, only the distribution of witness statistics impacts the structure and stability of the chain, assuming no simultaneous broadcasts, via mining events and stakeholder validation. Understanding this we were able to accelerate mining (Appendix D.2), and simulate quantum hashes by bootstrapping (Section D.3). We first describe our main statistical measures of blockchain stability in Section D.1.